what is the length of the major axis of the ellipse
3x(to the second power) + 4y(to the second power)=12

Found 2 solutions by stanbon, dolly:
Answer by stanbon(48564)

what is the length of the major axis of the ellipse
3x(to the second power) + 4y(to the second power)=12
--------------
Divide thru by 12 to get:
(x^2)/4 + (y^2)/3 = 1
a^2 = 4
a = 2
2a = 4
The length of the major axis is 2a = 4
===============
Cheers,
Stan H.

Answer by dolly(163)

Given equation of the ellipse is 3x^2 + 4y^2 = 12
Dividing by 12 throughout,
3x^2/12 + 4y^2/12 = 12/12
==> x^2/4 + y^2/3 = 1
This is of the form x^2/a^2 + y^2/b^2 = 1
On comparing we get a = 2
So length of major axis = 2a
= 4